OpenStudy (anonymous):

Find 2 numbers whose sum is 64 and whose difference is 42. 9 and 55 10 and 54 11 and 53 12 and 52

OpenStudy (nowhereman):

Why don't you just calculate the differences of those pairs of numbers?

OpenStudy (anonymous):

11 and 53

OpenStudy (anonymous):

X + Y = 64 .... 1 X - Y =42 .....2 ADD 1 AND 2 2X = 106 X = 53 BY EQUATION 1 X + Y =64 53 + Y = 64 Y = 11

OpenStudy (anonymous):

thanks LOVES (: